# Naturalistic Computational Cognitive Science Towards generalizable models and theories that capture the full range of natural behavior

### v1

@@ 4.2 Learning from naturalistic data can improve generalization
Studying how systems generalize is a key method in cognitive science and AI. Conceptual arguments often rely on empirical studies of generalization; however, these generalization properties are often studied in simple settings, with minimal models learning from minimal features (e.g. Marcus,, 1998). A fundamental assumption underlying this approach is that the simplifications do not alter the generalization problem. However, here we review studies from AI and neuroscience showing that the variability of experience according to seemingly-orthogonal variables can alter how systems learn and generalize—thus implying that studying models within more restricted settings may be misleading about the naturalistic computational problem.
@@
Example 1: Compositional generalization of image classifiers vs. agents. Naturalistic tasks can fundamentally change what models learn and how they generalize. Hill et al., (2020) trained two models to do a vision-language grounding task, and tested their compositional generalization to held-out language instructions. Both models were trained on the same language examples, and tested on the same held-out examples. However, one model was trained as an agent that interacts with a simulated environment, while the other was a simple image-language classifier trained on screenshots from that environment. The authors found that the interacting agent exhibited perfect compositional generalization to the novel language utterances, while the image classifier was substantially worse in novel settings. The authors also explored a range of other settings, including generalization benefits of more naturalistic 3D (rather than 2D) environments, or egocentric embodiments. Their results illustrate how richer, more naturalistic settings can enhance the generalizability of the solutions a system learns. Thus, if we are interested in understanding how a system generalizes, we may need to build models that learn from appropriately naturalistic data.
@@
Example 2: Generalization to novel syntactic structures is enhanced by variability in other structures’ fillers Generalization is of deep interest in linguistics. Recently, Misra and Mahowald, (2024) performed controlled experiments in which challenging linguistic constructions are systematically held out from the language model training data—and showed that models trained on naturalistic language can generalize to held out constructions by composing pieces of simpler constructions. Critically, the authors also found that generalization depended on the variability of the semantic fillers observed in the structures in training—thus showing how incorporating naturalistic variation in model training data can impact our theoretical inferences about generalization.
@@
Example 3: Rats raised in enriched environments. Analogously, animals raised in more complex environments can be more skilled than those raised in simpler environments. This phenomenon has been well-documented in neuroscience, where rats raised in enriched environments — with social interaction and/or more space or toys in their home cage — show greater exploration, and improved learning and memory on novel tasks (Simpson and Kelly,, 2011). This illustrates how, for natural as well as artificial intelligence, increasing the naturalistic variation in experience can alter the system’s learning in ways that may impact our experimental and theoretical inferences.
@@
One interesting finding from the deep learning literature is that when deep learning architectures with many parameters are trained with a lot of naturalistic data and an appropriate “basic” learning objective, these architectures can develop mechanisms that go beyond the learning objective they were trained on. Crucially, these learning paradigms can allow models to transfer well (through initial performance or accelerated learning) on novel tasks beyond the training distribution.
@@
Human learning may likewise benefit from transfer among the disparate tasks we learn. We argue that these AI findings motivate computational cognitive science research studying whether “down-stream human behavior” on a set of tasks can be recapitulated by a model which is trained on a large set of naturalistic tasks or stimuli representative of some of aspect of human experience—as in meta-learning approaches (cf. Wang,, 2021). Below, we provide examples of this kind of transfer spanning computer vision, reinforcement learning, and natural language processing.
@@
Example 1: Computer Vision. One of the earliest successes from deep learning came in computer vision. Soon after AlexNet (Krizhevsky et al.,, 2012) achieved strong results on the ImageNet dataset (Deng et al.,, 2009), researchers showed that the features discovered by AlexNet could be repurposed to novel tasks ranging from scene recognition to medical diagnosis (Donahue et al.,, 2014; Sharif Razavian et al.,, 2014; Litjens et al.,, 2017). This was striking because these features were trained (a) via supervision on a fixed set of objects but (b) enabled transfer to novel objects or even different types of tasks. For years this pattern continued, and researchers even showed that the representations learned by these models could (by 2022, unsurprisingly) transfer from Imagenet to both 3D dexterious manipulation tasks and 3D household navigation tasks, sometimes achieving better performance than using “ground-truth” state information (Parisi et al.,, 2022; Yuan et al.,, 2022). In addition to generally useful representations, deep learning vision models trained with naturalistic data also develop mechanisms they were not trained for. For example, scene-oriented CNNs develop mechanisms for object detection (Zhou et al.,, 2014) and vision transformers develop mechanisms for segmenting objects (Caron et al.,, 2021)—both without an explicit training signal.
@@
Example 2: Reinforcement learning. Likewise in reinforcement learning, researchers have found that reinforcement learning algorithms trained on many (e.g. billions) of tasks exhibit strong “out-of-distribution” generalization on unknown tasks (Team et al.,, 2021), with the ability to generalize to novel tasks in the same number of samples as humans (Team et al.,, 2023). Some algorithms can even generalize to collaborating with humans without any human data (Strouse et al.,, 2021). Recently, reinforcement learning algorithms for learning a modern variant of the successor representation (Gershman,, 2018; Carvalho et al.,, 2024) have shown that they develop mechanism for exploration and behavioral skills without explicit training signals (Liu et al.,, 2024).
@@
Example 3: Large language models. Perhaps the most striking example of this general phenomenon comes from large language models. Large language models are only trained to predict the next “token” (or word) to appear following a sequence of tokens. However, the distribution of internet language effectively includes a broad mixture of many tasks (Radford et al.,, 2019). When trained on vast amounts of this naturalistic data, large language models develop mechanisms for disparate tasks such as modular arithithmic, solving word problems, etc. (Wei et al.,, 2022)—and even for adapting to new tasks from examples presented in context (Brown et al.,, 2020). Moreover, the representations learned by these models can transfer to many superficially dissimilar downstream tasks (Lu et al., 2022a, ), even ones as dissimilar as playing video games (Reid et al.,, 2022). These examples illustrate how learning from a broad naturalistic distribution can induce many abilities—and can provide important transfer to downstream tasks.
@@ 4.4 Learning from naturalistic data allows us to ask new questions about the origins of knowledge
The points outlined above have an important consequences; using models that learn from naturalistic data can change our theoretical conclusions in cognitive science. For example, if a model fails to generalize when trained on simple data, we cannot be sure if the model or the data are inadequate. By contrast, using models that learn from naturalistic data can enable us to ask more precise questions about which features—of the models or the data—are necessary to reproduce the theoretically-relevant features of cognitive and neural processes.
@@
Example 1: language models and the learnability of language. Recent progress in language modeling from naturalistic data has suggested challenges to prior assumptions about language learnability and innateness. We briefly sketch these developments; Piantadosi, (2023) and Futrell and Mahowald, (2025) give much more thorough accounts. Classical approaches to language processing focused on simple models and ideas (e.g. Chomsky,, 2014), but in doing so imposed strong assumptions. Under certain strong assumptions, Gold, (1967) proved that even relatively simple languages cannot be learned from input alone. This proof contributed to arguments that there must be some innate universals such as recursion, and separation of syntax from semantics (Chomsky,, 1957, 1965, e.g.)—ideas that were subsequently influential in arguments about the nature of cognition more broadly (Fodor et al.,, 1975; Fodor and Pylyshyn,, 1988). These theories were based on studying particular language features in great detail [^2] Or sometimes even studying the intuitions of linguists, with which the average language user would not necessarily even agree (Spencer,, 1973).. Because of this focus, the theories did not attempt to model language processing in all its naturalistic detail.
@@
However, recent progress in deep learning based language models (e.g. Brown et al.,, 2020) shows that many of these assumptions may need revision. By learning from large amounts of naturalistic data, these models acquire both syntax (Manning et al.,, 2020) and semantics (Li et al., 2021a, )—and account for a broad range of behavioral and neuroscientific phenomena (Schrimpf et al.,, 2021). Of course, these models often learn from inhuman quantities of language (Wilcox et al.,, 2024). However, even language models trained on a human-like amount of language data can learn complex syntactic features that have been previously considered evidence for innate gramatical knowledge (e.g. Wilcox et al.,, 2023). Likewise, while some “impossible” languages that (adult) humans struggle to learn had been considered evidence for universal grammar, Kallini et al., (2024) show that language models can more easily learn real languages than impossible ones. Finally, controlled experiments—like those from Misra and Mahowald, (2024), reviewed above—allow us to begin to understand which features of experience can contribute to generalizing to truly-novel structures, and how naturalistic variability contributes to this generalization. Thus, incorporating naturalistic data into has helped to reshape our understanding of the necessary and sufficient features for acquiring linguistic capabilities.
@@
Example 2: naturalistic data can lead to brain-like functional specialization. Given the importance of perceiving faces in our social lives, it might seem likely that face recognition is innately encoded in the brain. Indeed, the discovery of the Fusiform Face Area (FFA)—a visual region specialized for face perception (Kanwisher and Yovel,, 2006)—would seem to lend further support to this hypothesis. However, recent computational work shows that this kind of specialization can emerge from a domain-general computer-vision model purely through training on a naturalistic distribution of object and face recognition tasks (Dobs et al.,, 2022). As the authors state: “It may be that the only inductive bias humans need to develop their face system is the already well-established early preference of infants to look at faces.” Of course, this does not necessarily imply that the FFA is not innately specified. However, it does illustrate how training computational models over naturalistic data can help us to understand the natural constraints that affect neural organization—whether those effects occur over developmental or evolutionary timescales.
@@ 5 Building generalizable models
ML research has excelled in developing generalizable models. While we focus on their utility for working on naturalistic stimuli and tasks, we believe cognitive science would broadly benefit from improved practices for developing generalizable models. In this section, we highlight three key practices of empirical ML research that we believe computational cognitive science can adopt. The first is frictionless reproducibility: the practice of developing research artifacts which can be reused, and repurposed with minimal effort (§5.1). The second is a focus on developing generalizable models that can learn successfully across many datasets (§5.2). We argue that these two offer the foundation for having groups of research collectively work on important research problems through the utilization of benchmarks (§5.3).
Before we continue, we acknowledge that modern empirical ML is a young field which, like other fields, has faced challenges in rigor and reproducibility (Melis et al.,, 2017; Lucic et al.,, 2018; Riquelme et al.,, 2018; Henderson et al.,, 2018; Recht et al.,, 2019; Agarwal et al.,, 2021). Despite these challenges, we are optimistic that some of the practical knowledge ML has developed is valuable for improving rigor and reproducibility, as we highlight below.
Figure 6: Overview of key strategies cognitive science can adopt to develop generalizable models. (Figures reproduced from (Geirhos et al.,, 2018; Bowers et al.,, 2023))
Section summary. We envision a future where cognitive scientists identify a cognitive phenomenon of interest (e.g., theory of mind, working memory, or object recognition) and create a phenomena-oriented benchmark (R8) containing experiments specifically designed to test different theoretical aspects of that phenomenon. Here, experiments are implemented with standardized evaluation protocols and interfaces (R1) that enable direct comparison between competing theoretical frameworks. All experimental data and stimuli are made available through one-line download access (R2). Models are developed and released following frictionless reproducibility standards (R1-R3) for easy adoption and modification. Researchers identify gaps in theoretical coverage and contribute new experiments to the benchmark, facilitating maintaining a separation between model and task design (R6). Thanks to the existence of a benchmark and frictionless reproducibility standards, models are evaluated comprehensively across all experimental paradigms with minimal additional human labor. This helps mitigate the challenge of disconnected, incommensurate findings that resist integration across studies (Almaatouq et al.,, 2024). Thanks to the leaderboard principle (R7), this also helps ensure that progress is defined by substantial rather than incremental advances. This approach encourages the development of models that work across both simplified and naturalistic contexts (R4) and that generalize across multiple experimental paradigms (R5). Importantly, in addition to a model that predicts behavior across many tasks, we also have a model that can perform many tasks that humans can—enabling the building of human-like models that scale and generalize.
@@ 5.1 Frictionless reproducibility: a superpower
In recent years, data-centric sciences are experiencing a profound transition to what David Donoho calls “frictionless reproducibility” (Donoho,, 2024; cf. Recht,, 2024). Donoho identifies three key pillars to frictionless reproducibility: data sharing, code sharing, and competitive challenges. We will focus on the first two in this section, and competitive challenges in §5.3. Critically, with frictionless reproducibility, the research artifacts that are produced allow future researchers to exactly re-execute the same complete workflow with minimal effort. With this, good ideas are spread, adopted, and improved with little friction. Donoho argues that ML is ostensibly the most successful at frictionless reproducibility. Below we review some supporting strategies that may aid cognitive science.
@@
R1. Use a standardized evaluation protocol where models can be compared via a common interface. Machine learning’s foundation of standardized evaluation protocols traces back to 1959, when Highleyman created the first alphanumeric dataset and the first ever “train-test” split for comparing pattern recognition (Highleyman and Kamentsky,, 1959). This strategy continued with the famous MNIST alphanumeric dataset (LeCun et al.,, 1998). MNIST simplified data-sharing and evaluation by pre-processing data, fixing the train and test distributions, and sharing the data online. This standardization and usage of a common interface enabled fair comparisons between diverse approaches, from boosted stumps (Kégl and Busa-Fekete,, 2009) to support vector machines (Cristianini and Scholkopf,, 2002) to nearest neighbor methods (Khan,, 2017).
Standardized evaluation protocols serve a critical function: mitigating personal bias in model assessment. The evolution from MNIST through Caltech 101 (Fei-Fei et al.,, 2004) to ImageNet (Deng et al.,, 2009) demonstrates this principle, as increasingly naturalistic datasets enabled fair comparisons between diverse model families. This enabled AlexNet (Krizhevsky et al.,, 2012) to validate deep learning in an era when many researchers doubted its potential.
Thus, we recommend that cognitive scientists establish clear hold-out data for either model development or data analysis. This is especially critical for confirming exploratory results (Poldrack et al.,, 2017). Combined with standardized evaluation protocols, this will enable fair comparison between diverse theoretical frameworks ranging from program synthesis to deep neural networks to reinforcement learning and active inference.
R2. Make data accessible with a single line of code. Fully processed research data should be programmatically accessible with a single line of code, requiring no permissions. While uploading data to platforms like the Open Science Framework (OSF) is valuable, we additionally recommend uploading data to platforms like Hugging Face that offer version control, free hosting, authentication-free access, the ability to download data with one line of code, and to upload data with a few lines of code. This removes friction from reusing and updating datasets.
@@
R3. Adopt the “single file” philosophy. Ideally, key workflows (e.g. model or evaluation definititions) should be defined in a single file. This has emerged as a powerful force for reproducibility in ML research (see Table 1 for examples). While this approach contradicts traditional software engineering principles of modularity and abstraction, it offers distinct advantages for research code: (1) complete understanding of an important component (e.g. how a model works) requires reading only one file, (2) minimal abstraction layers and dependencies, and (3) easy adaptation of either complete implementations or specific components. One can simply copy this single file into another codebase. Removing friction from future reuse also benefits us—as we can see from Table 1, these methods get a large number of citations.

### v2

@@ 4.2 Learning from naturalistic data can improve generalization
Studying how systems generalize is fundamental to cognitive science and AI. When studying generalization in simplified settings, a key assumption is that the simplifications do not alter the generalization problem. Here, we review studies from AI and neuroscience showing that variability in experience (even aloong seemingly-orthogonal axes) can alter learning and generalization—thus implying that studying models within more restricted settings may mislead us about the naturalistic computational problem.
@@
Example 1: Compositional generalization of image classifiers vs. agents. Naturalistic tasks can fundamentally change what models learn and how they generalize. Hill et al., (2020) trained two models on a vision-language grounding task, and tested their compositional generalization to held-out instructions. Both models were trained on the same language examples, and tested on the same held-out examples. However, one model was trained as an agent interacting with a simulated environment, while the other was simply trained trained to classify screenshots from that environment based on language instructions. The authors found that the interacting agent exhibited perfect compositional generalization to novel language utterances, while the image classifier generalized substantially worse. The authors also explored other factors, including generalization benefits of more naturalistic 3D (rather than 2D) environments, or egocentric embodiments. Their results illustrate how richer naturalistic settings can enhance the generalizability of learned solution. Thus, if we are interested in understanding generalization, we may need to build models that learn from appropriately naturalistic data.
@@
Example 2: Syntactic generalization is enhanced by variability in other structures’ fillers Structural generalization is of deep interest in linguistics. Recently, Misra and Mahowald, (2024) performed controlled experiments in which challenging linguistic constructions are systematically held out from language model training data, and showed that models trained on naturalistic language generalize to held out constructions by composing pieces of simpler constructions. Critically, this generalization depended on sufficient variability in the semantic fillers observed in the structures seen in training—thus, incorporating naturalistic variation in model training data can impact our theoretical inferences about generalization.
@@
Example 3: Rats raised in enriched environments. Analogously, neuroscientists have found that animals raised in more complex environments can be more skilled than those raised in simpler environments. Rats raised in enriched environments — with social interaction and/or more space or toys — show greater exploration, and improved learning and memory on novel tasks (Simpson and Kelly,, 2011). Thus, for natural as well as artificial intelligence, greater naturalistic variation in experience can alter the system’s learning in ways that impact our experimental and theoretical inferences.
@@
One interesting finding from AI is that when highly-parameterized deep learning architectures are trained with bountiful naturalistic data and appropriate “basic” learning objectives, these architectures can develop mechanisms that go beyond the training objective. Crucially, these learning paradigms can allow models to transfer well (through initial performance or accelerated learning) on novel tasks beyond the training distribution.
@@
Human learning may likewise benefit from transfer among the disparate tasks we learn. We argue that these AI findings should motivate cognitive science research studying whether “downstream human behavior” on a set of tasks can be recapitulated by a model which is trained on varied naturalistic tasks or stimuli representative of (some aspects of) human experience. Below, we provide examples illustrating this kind of transfer.
@@
Example 1: Computer Vision. One early success of deep learning came in computer vision. Soon after AlexNet (Krizhevsky et al.,, 2012) achieved strong results on the ImageNet dataset (Deng et al.,, 2009), researchers showed that AlexNet’s feature could be repurposed to novel tasks like scene recognition or medical diagnosis (Donahue et al.,, 2014; Sharif Razavian et al.,, 2014; Litjens et al.,, 2017). This result was striking because features trained via supervision on fixed object classes nevertheless enabled transfer to novel objects or even different types of tasks. This pattern has continuend, and more recently researchers showed that the representations learned by these models could transfer from Imagenet 3D household navigation and dexterous tasks, sometimes achieving better performance than using “ground-truth” state (Parisi et al.,, 2022; Yuan et al.,, 2022). Beyond their useful representations, vision models trained with naturalistic data also develop mechanisms they were not trained for. For example, scene-oriented CNNs develop mechanisms for object detection (Zhou et al.,, 2014) and vision transformers develop mechanisms for segmentation (Caron et al.,, 2021)—both without an explicit training signal.
@@
Example 2: Reinforcement learning. Likewise in reinforcement learning, researchers have found that agents trained on many (billions) of tasks exhibit strong “out-of-distribution” generalization on unknown tasks (Team et al.,, 2021), learning novel tasks as efficiently as humans (Team et al.,, 2023). Some algorithms can even generalize to collaborating with humans without any human data (Strouse et al.,, 2021). Recently, reinforcement learning algorithms for learning a modern variant of the successor representation (Gershman,, 2018; Carvalho et al.,, 2024) have been shown to develop skills for exploration and behavior without explicit training signals (Liu et al.,, 2024).
@@
Example 3: Large language models. Perhaps the most striking example of this general phenomenon comes from large language models. These models are only trained to predict the next “token” (or word) to appear in a sequence. However, the distribution of internet language effectively includes a broad mixture of many tasks (Radford et al.,, 2019). When trained on these vast naturalistic data, language models develop mechanisms for disparate tasks such as modular arithithmic, solving word problems, etc. (Wei et al.,, 2022)—and even for adapting to new tasks from examples (Brown et al.,, 2020). Moreover, the representations learned by these models can transfer to superficially dissimilar downstream tasks (Lu et al.,, 2022), including playing video games (Reid et al.,, 2022). These examples illustrate how learning from a broad naturalistic distribution can induce many abilities—and can transfer to downstream tasks.
@@ 4.4 Learning from naturalistic data allows us to ask new questions about the origins of knowledge
The points outlined above have important consequences: using models that learn from naturalistic data can change our theoretical conclusions in cognitive science. For example, if a model fails to generalize from simple data, we cannot be sure whether the model or the data are inadequate. By contrast, using models that learn from naturalistic data can enable us to ask more precise questions about which features—of models or data—are necessary to reproduce the theoretically-relevant features of cognitive and neural processes.
@@
Example 1: language models and learnability. Recent progress in language modeling has suggested challenges to prior assumptions about language learnability and innateness. We briefly sketch these issues; see Piantadosi, (2023) and Futrell and Mahowald, (2025) for thorough accounts. Classical approaches to language processing focused on simple models and ideas (e.g. Chomsky,, 2014), but in doing so imposed strong assumptions. Under certain strong assumptions, Gold, (1967) proved that even relatively simple languages cannot be learned from input alone. This proof contributed to arguments that for innate universals such as recursion, and separation of syntax from semantics (Chomsky,, 1957, 1965, e.g.)—ideas that were subsequently influential in arguments about the nature of cognition more broadly (Fodor et al.,, 1975; Fodor and Pylyshyn,, 1988; Quilty-Dunn et al.,, 2023). However, these theories did not attempt to model language processing in all its naturalistic detail.
@@
By contrast, when deep-learning models learn from large amounts of naturalistic data, these models acquire aspects of both syntax (Manning et al.,, 2020) and semantics (Li et al., 2021a, )—and account for many behavioral and neuroscientific phenomena (Schrimpf et al.,, 2021). Although these models often learn from inhuman quantities of language (Wilcox et al.,, 2024), even models trained on a human-like quantity can learn complex syntactic features that were previously considered evidence for innate knowledge (e.g. Wilcox et al.,, 2023). Likewise, while some “impossible” languages that (adult) humans struggle to learn had been considered evidence for universal grammar, Kallini et al., (2024) show that language models can more easily learn real languages than impossible ones. Finally, controlled experiments—like those from Misra and Mahowald, (2024), reviewed above—begin to illustrate which features of experience contribute to syntactic generalization: particularly the importance of naturalistic variation. Thus, studying learnability from naturalistic data has helped to reshape our understanding of the necessary and sufficient features for acquiring linguistic capabilities.
@@
Example 2: naturalistic data can lead to brain-like functional specialization. Given the social importance of face perception, it might seem likely that face recognition is innately encoded in the brain. Indeed, the discovery of the Fusiform Face Area (FFA)—a visual region specialized for face perception (Kanwisher and Yovel,, 2006)—supports this hypothesis. However, recent computational work shows that this kind of specialization can emerge from a domain-general computer-vision model purely through training on a naturalistic distribution of object and face recognition tasks (Dobs et al.,, 2022). As the authors state: “It may be that the only inductive bias humans need to develop their face system is the already well-established early preference of infants to look at faces.” Of course, this does not necessarily imply that the FFA is not innately specified. However, it does illustrate how training computational models over naturalistic data can help us generate hypotheses about the causes of our neural organization.
@@ 5 Building generalizable models
Figure 6: Overview of key strategies cognitive science can adopt to develop generalizable models. (Figures reproduced from Geirhos et al.,, 2018; Bowers et al.,, 2023)
ML research focuses on developing generalizable models. While we focus on their utility for working with naturalistic tasks, we believe cognitive science can broadly benefit from improved practices for developing generalizable models. Here, we highlight three key practices of empirical ML research that we believe computational cognitive science can adopt. The first is frictionless reproducibility: developing research artifacts that can be reused and repurposed with minimal effort (§5.1). The second is a focus on developing generalizable models that can learn successfully across many datasets (§5.2). We argue that these two offer the foundation for the third: collectively collaborating on important research problems through dynamic benchmarks (§5.3).
Before we continue, we acknowledge that modern empirical ML is a young field which, like other fields, has faced challenges in rigor and reproducibility (Melis et al.,, 2017; Lucic et al.,, 2018; Riquelme et al.,, 2018; Henderson et al.,, 2018; Recht et al.,, 2019; Agarwal et al.,, 2021). Despite these challenges, we believe that some of the practices of ML are valuable for improving rigor and reproducibility, as we highlight below.
@@ 5.1 Frictionless reproducibility: a superpower
In recent years, data-centric sciences are experiencing a profound transition to what David Donoho calls “frictionless reproducibility” (Donoho,, 2024; cf. Recht,, 2024). Donoho identifies three key pillars to frictionless reproducibility: data sharing, code sharing, and competitive challenges. We focus on the first two in this section, and competitive challenges in §5.3. Critically, with frictionless reproducibility, the research artifacts enable future researchers to exactly re-execute the complete workflow with minimal effort. Thus, good ideas are spread, adopted, and improved with little friction. Donoho argues that ML is ostensibly the most successful at frictionless reproducibility. Below we review some strategies that may aid cognitive science.
@@
R1. Use standardized evaluation protocols where models are compared via a common interface. Machine learning’s foundation of standardized evaluation protocols traces back to 1959, when Highleyman and Kamentsky, (1959) created the first alphanumeric pattern-recognition dataset and the first ever “train-test” split. This strategy continued with the famous MNIST dataset (LeCun et al.,, 1998). MNIST simplified data-sharing and evaluation by pre-processing data, fixing train and test distributions, and sharing the data online. This standardization enabled fair comparisons between diverse approaches, from boosted stumps (Kégl and Busa-Fekete,, 2009) to support vector machines (Cristianini and Scholkopf,, 2002) or nearest neighbors (Khan,, 2017).
Standardized evaluation protocols serve a critical function: mitigating personal bias in model assessment. The evolution to Caltech 101 (Fei-Fei et al.,, 2004) and ImageNet (Deng et al.,, 2009) demonstrates this principle, as increasingly naturalistic datasets enabled fair comparisons between diverse model families. This enabled AlexNet (Krizhevsky et al.,, 2012) to validate deep learning in an era when many researchers doubted its potential.
Thus, we recommend that cognitive scientists establish clear hold-out data for both model development and data analysis. This is especially critical for confirming exploratory results (Poldrack et al.,, 2017). Combined with standardized evaluation protocols, this strategy enables fair comparison between diverse theoretical frameworks ranging from program synthesis to deep neural networks to reinforcement learning.
R2. Make data accessible with a single line of code. Fully processed research data should be programmatically accessible with a single line of code, requiring no permissions. While uploading data to platforms like the Open Science Framework (OSF) is valuable, we additionally recommend platforms like Hugging Face that offer version control, free hosting, authentication-free access, access with one line of code, and uploading data with a few lines of code. This removes friction from reusing and updating datasets.
@@
R3. Adopt the “single file” philosophy. Ideally, key workflows (e.g. model or evaluation definitions) should be defined in a single file. This has emerged as a powerful force for reproducibility in ML. While this approach contradicts traditional software engineering principles of modularity and abstraction, it offers distinct advantages for research code: (1) complete understanding of an important component (e.g. a modell) requires reading only one file, (2) minimal abstraction layers and dependencies, and (3) easy adaptation of prior methods or their components, by simply copying this single file into another codebase. Removing friction from future reuse also benefits us—as Table 1 shows, single-file methods get many citations.